# AgenticMath: Enhancing LLM Reasoning via Agentic-based Math Data Generation

### v2

@@ Abstract
The creation of high-quality datasets to improve Large Language Model (LLM) reasoning remains a significant challenge, as current methods often suffer from generating low-quality/incorrect answers and limited information richness from available data sources. To address this, we propose AgenticMath, a novel agentic pipeline for generating high-quality mathematical question-answer pairs to enhance the supervised fine-tuning of LLMs. Our method operates through four stages: (1) Seed Question Filter that selects questions with high information richness, complexity, and clarity; (2) an Agentic Question Rephrase step that employs a multi-agent system to generate diverse, logically consistent paraphrases; (3) an Answer Augment step where rewrite answers using chain-of-thought reasoning to enhance numerical and logical correctness, without reliance on human-provided labels; and (4) a final Question and Answer Evaluation that retains only the most superior pairs. Extensive experiments demonstrate that, fine-tuning 3B-8B parameter LLMs on AgenticMath generated datasets (comprising only 30-60K math samples) achieves competitive or superior performance on diverse in domain and out-of-domain mathematical reasoning benchmarks compared to baselines trained on much more data (e.g., 400K or 2.3M samples). Our work demonstrates that targeted, high-quality data generation is a more efficient path to improving mathematical reasoning in LLMs than large-scale, low-quality alternatives.
@@ 2 Related Work
Large language models (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023a, b; Chowdhery et al., 2023; Bi et al., 2024; Team et al., 2023, 2024; Grattafiori et al., 2024) show strong general ability and are increasingly applied to mathematical problem solving (Cobbe et al., 2021; Hendrycks et al., 2021; Zhang et al., 2024a; Xia et al., 2025). Prompt-based approaches (Wei et al., 2022; Wang et al., 2022; Fu et al., 2022) extend reasoning paths but yield limited improvements. Recent work thus emphasizes synthesizing math reasoning data for supervised fine-tuning (Yu et al., 2023; Luo et al., 2023; Tang et al., 2024; Li et al., 2024a; Zhang et al., 2024b; Liu et al., 2025a; Tong et al., 2024). WizardMath (Luo et al., 2023) adds evolution directives and reinforcement learning; MathFusion (Pei et al., 2025) fuses problems for relational reasoning. Other methods integrate code tools (Yue et al., 2023; Wang et al., 2023; Hosseini et al., 2014; Toshniwal et al., 2024; Li et al., 2024b; Lu et al., 2024). In this work, we advance mathematical reasoning by improving both the question formulation and answer quality in synthetic data.
Multi-Agent for Data Generation.
Multi-agent systems (Hong et al., 2023; Wu et al., 2023; Li et al., 2023; aut, 2023) show strong ability and are increasingly applied to data synthesis. Abdullin et al. (2024) proposed a multi-intelligence framework for dialog generation, while MAGDi (Chen et al., 2024a) used graph-based interactions and MALLM-GAN (Ling et al., 2024) employed generator–discriminator agents for tabular data. AgentInstruct (Mitra et al., 2024a) and Orca-Math (Mitra et al., 2024b) iteratively refined instructions, whereas role-driven synthesis was explored by Ge et al. (2024) and VCR (Liu et al., 2025b). MALT (Motwani et al., 2024) introduced generator, verifier, and refiner agents for math problems. Despite these advances, ensuring high-quality data for mathematical reasoning remains challenging. Hence, we introduce seed filtering and quality evaluation agents to ensure reliable math reasoning data.
Data Selection.
Data selection has been shown to improve instruction tuning, achieving performance comparable to full training datasets (Xia et al., 2022; Lu et al., 2023; Xia et al., 2023; Li et al., 2024c; Xia et al., 2024; Pang et al., 2024; Zhang et al., 2024c; Long et al., 2024; Li et al., 2025a; Wang et al., 2025; Xu et al., 2025; Pang et al., 2025; Li et al., 2025b). Specifically, LESS (Pang et al., 2024) proposes introduces a low-rank gradient similarity–based selection method. DS2 (Pang et al., 2024) further refines data quality by rating and curating samples with LLM. In this work, we integrate LLM-based selection into data generation, allowing synthesis conditioned solely on high-quality questions.
@@ 3 AgenticMath: Multi-Agent Design for Math Reasoning
This section details the proposed AgenticMath(see LABEL:fig1), which design Agentic workflow to generate high quality math problems and reasoning solutions based on seed data GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). The framework combined with four stages: Seed problem filtering, Agentic problem generation, Reasoning solution generation, and Synthetic data evaluation. We generate a high quality math dataset to finetune LLMs enhancing their math reasoning ability by using AgenticMath. All prompts used for the agents are provided in Appendix A.5.
@@ 3.2 Stage 1: Seed Problem Filtering
Figure 1: Workflow of Stage 1 showing the filtering process that removes low-quality seed problems and retains high-value ones for subsequent synthesis.
@@
Problem Filtering Dimensions and Prompt Example Complexity: Does it integrate multiple mathematical domains (e.g., algebra + geometry) or demand critical thinking? Information Value: Does it contain useful knowledge or reasoning opportunities? Can it help learners discover concepts, strategies, or patterns? Clarity & Precision: Is the question unambiguous, logically consistent, and free of errors? Poorly framed questions score lower.
@@ 3.3 Stage 2: Agentic Problem Generation
Figure 2: Workflow of Stage 2 showing multi-agent problem generation with Rephrase, Review, and Revise agents, along with iterative refinement to ensure clarity, coherence, and mathematical validity.
@@
Problem Review Dimensions and Prompt Example Clarity & Grammar: The question must be grammatically correct, precisely phrased, and easy to understand. It should avoid ambiguity in wording or phrasing. Logical Coherence & Completeness: All elements of the problem (e.g., given information, constraints, relationships, objectives) must be logically interconnected and sufficient. The problem should present a clear, sequential path for reasoning, without missing information required for the specified solution approach. Mathematical Validity & Solvability: The problem must be fundamentally a mathematics problem, with all its premises and conditions being mutually consistent and mathematically sound…
@@ 3.4 Stage 3: Solution Generation
Figure 3: Workflow of Stage 3 showing solution generation by the Solver Agent with step-by-step reasoning.
@@
$\mathcal{D}^{\text{sol}}_{\text{seed}}=\{(q_{i}^{\text{seed}},\,a_{i}^{\text{sol}})\}_{i=1}^{N},\quad\mathcal{D}^{\text{sol}}_{\text{ref}}=\{(q_{j}^{\text{ref}},\,a_{j}^{\text{sol}})\}_{j=1}^{K},$
@@ 3.5 Stage 4: Synthetic Data Evaluation
Figure 4: Workflow of Stage 4 showing evaluation of problem–solution pairs for quality and diversity.
In this stage, the scoring and curation framework from Stage 1 is extended to problem–solution pairs. The evaluation targets the synthetic set $\mathcal{D}^{\text{sol}}_{\text{ref}}=\{(q_{j}^{\text{ref}},a_{j}^{\text{sol}})\}_{j=1}^{K}$. Each pair is judged along three dimensions—clarity of the problem, correctness of the solution, and completeness of reasoning. Scores are further stabilized using the Score Transition Matrix and refined through $k$-NN consistency checks. To build a high-quality and diverse subset, a ranking-based selection is applied instead of a fixed threshold. Pairs are first sorted by quality and grouped into discrete score levels. Within each group, ranking is refined by long-tail diversity, favoring samples that are more distant in the embedding space and thus add unique information. Selection proceeds from high to low quality, with diversity guiding intra-group choice, until L=15, [^0] L=15{,}000 pairs are collected. This yields the curated dataset $\mathcal{D}_{\text{selected}}^{\text{sol}}=\{(q_{j}^{\text{ref}},a_{j}^{\text{sol}})\}_{j=1}^{L}$. The final training dataset combines this curated set with the seed-based solutions: $\mathcal{D}_{\text{final}}=\mathcal{D}_{\text{selected}}^{\text{sol}}\cup\mathcal{D}^{\text{sol}}_{\text{seed}}$, ensuring both rigor and diversity for downstream fine-tuning.
@@
DeepSeekMath-7B-MathFusion 195K $58.2$ $79.5$ $40.3$ $69.1$ $25.5$ $27.0$ $49.9$
AgenticMath-DSMath-7B 30K $52.4$ $80.1$ $42.6$ $66.8$ $18.2$ $26.9$ $47.8$
AgenticMath-DSMath-7B 60K $55.0$ $80.1$ $43.6$ $69.9$ $20.0$ $27.0$ $49.3$
@@
AgenticMath-Llama3-8B 30K $36.8$ $78.4$ $29.6$ $40.3$ $11.4$ $20.4$ $36.2$
AgenticMath-Llama3-8B 60K $40.4$ $80.1$ $31.6$ $46.7$ $14.1$ $22.6$ $39.3$
AgenticMath-Llama3-8B 90K $42.8$ $81.4$ $33.0$ $48.6$ $13.9$ $21.8$ $40.3$
@@ 4.1 Experimental Setup
We compare AgenticMath with state-of-the-art methods in two settings. For large-scale training, we include MetaMath (Yu et al., 2023), RFT (Yuan et al., 2023), DART-Math (Tong et al., 2024), MathScale (Tang et al., 2024), DeepSeekMath-7B-Instruct (sha, 2024), RefAug (Zhang et al., 2024b), MMIQC (Liu et al., 2025a), and WizardMath (Luo et al., 2023), all using 400K–2.3M samples. For small-scale, we evaluate 30K and 60K subsets: RefAug and MathFusion provide native 30K versions, while other baselines are randomly down-sampled to 60K from the large-scale dataset above.
@@ 4.3 Understanding AgenticMath: Ablations and Insights
Table 3: Ablation study on the contribution of different pipeline stages.
Samples AVG
Mistral-7B: GSM8K + MATH 15K 18.0
   + Stage 3: Solution Augmentation 30K 34.7 $\uparrow 16.7$
   + Stage 1: Problem Filtering 29K 34.9 $\uparrow 0.2$
   + Stage 2: Problem Rephrased 30K 35.2 $\uparrow 0.3$
   + Stage 2: Problem Review-Revise Loop 30K 35.6 $\uparrow 0.4$
   + Stage 4: Synthetic Data Evaluation 30K 35.8 $\uparrow 0.2$
To better understand the contribution of each stage in the AgenticMath pipeline, we conduct ablation studies on Mistral-7B with a fixed training set size of 30K samples. Table 3 shows that adding solution augmentation (Stage 3) brings the largest single improvement, confirming the importance of rich reasoning traces. Problem filtering (Stage 1) and rephrasing (Stage 2) both provide further gains, while the review–revise loop improves the quality of rephrasings that initially fail to meet the threshold. The final step, joint problem–solution evaluation (Stage 4), yields the highest overall performance. Overall, these results highlight that AgenticMath achieves consistent improvements by enforcing quality control at every stage, making it more effective than one-shot data synthesis approaches.
@@
Larger Training Samples Yield Stronger Reasoning Performance.
@@
Illustrative Cases of Enhanced Problem Quality
Appendix A.4 provides several illustrative cases refined by the Reviewer and Revise Agents, showing how our method improves clarity and correctness of mathematical problems.
5 CONCLUSION
Appendix A Appendix
A.1 The Use of Large Language Models (LLMs)
A.2 Training Details

### v3

@@ Abstract
The creation of high-quality datasets to improve Large Language Model (LLM) reasoning remains a significant challenge, as current methods often suffer from generating low-quality/incorrect answers and limited information richness from available data sources. To address this, we propose AgenticMath, a novel agentic method for generating high-quality mathematical question-answer pairs to enhance the supervised fine-tuning of LLMs. Our method operates through four stages: (1) Seed Question Filter that selects questions with high information richness, complexity, and clarity; (2) an Agentic Question Rephrase step that employs a multi-agent system to generate diverse, logically consistent paraphrases; (3) an Answer Augment step where rewrite answers using chain-of-thought reasoning to enhance numerical and logical correctness, without reliance on human-provided labels; and (4) a final Question and Answer Evaluation that retains only the most superior pairs. Extensive experiments demonstrate that, fine-tuning 3B-8B parameter LLMs on AgenticMath generated datasets (comprising only 30-60K math samples) achieves competitive or superior performance on diverse in domain and out-of-domain mathematical reasoning benchmarks compared to baselines trained on much more data (e.g., 400K or 2.3M samples). Our work demonstrates that targeted, high-quality data generation is a more efficient path to improving mathematical reasoning in LLMs than large-scale, low-quality alternatives.
AgenticMath: Enhancing LLM Reasoning via
Agentic-based Math Data Generation
Xianyang Liu1   Yilin Liu3   Shuai Wang3   Hao Cheng2 Andrew Estornell4   Yuzhi Zhao5   Jun Shu6   Jiaheng Wei3thanks: Corresponding to: jiahengwei@hkust-gz.edu.cn 1King’s College London  2Hong Kong Baptist University 3The Hong Kong University of Science and Technology (Guangzhou)  4ByteDance Seed 5City University of Hong Kong  6Xi’an Jiaotong University liuxianyang98@gmail.comjiahengwei@hkust-gz.edu.cn
@@ 2 Related Work
Large language models (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023a, b; Chowdhery et al., 2023; Bi et al., 2024; Team et al., 2023, 2024; Grattafiori et al., 2024) show strong general ability and are increasingly applied to mathematical problem solving (Cobbe et al., 2021; Hendrycks et al., 2021; Zhang et al., 2024a; Xia et al., 2025). Prompt-based approaches (Wei et al., 2022; Wang et al., 2022; Fu et al., 2022) extend reasoning paths but yield limited improvements. Recent work thus emphasizes synthesizing math reasoning data for supervised fine-tuning (Yu et al., 2023; Luo et al., 2023; Tang et al., 2024; Li et al., 2024a; Zhang et al., 2024c; Liu et al., 2025a; Tong et al., 2024; Ding et al., 2025). WizardMath (Luo et al., 2023) adds evolution directives and reinforcement learning; MathFusion (Pei et al., 2025) fuses problems for relational reasoning. Other methods integrate code tools (Yue et al., 2023; Wang et al., 2023; Hosseini et al., 2014; Toshniwal et al., 2024; Li et al., 2024b; Lu et al., 2024). In this work, we advance mathematical reasoning by improving both the question formulation and answer quality in synthetic data.
Multi-Agent for Data Generation
Multi-agent systems (Hong et al., 2023; Wu et al., 2023; Li et al., 2023; aut, 2023) show strong ability and are increasingly applied to data synthesis. Abdullin et al. (2024) proposed a multi-intelligence framework for dialog generation, while MAGDi (Chen et al., 2024a) used graph-based interactions and MALLM-GAN (Ling et al., 2024) employed generator–discriminator agents for tabular data. AgentInstruct (Mitra et al., 2024a) and Orca-Math (Mitra et al., 2024b) iteratively refined instructions, whereas role-driven synthesis was explored by Ge et al. (2024) and VCR (Liu et al., 2025b). MALT (Motwani et al., 2024) introduced generator, verifier, and refiner agents for math problems. Despite these advances, ensuring high-quality data for mathematical reasoning remains challenging. To address this, we introduce seed filtering and quality evaluate agents to guarantee reliable math reasoning data.
@@ 3 AgenticMath: Multi-Agent Design for Math Reasoning
This section details the proposed AgenticMath (see Figure˜1), which is designed to generate high-quality mathematical problems and reasoning solutions based on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) seed datasets. The framework consists of four stages: seed problem filtering, agentic problem generation, reasoning-solution generation, and synthetic-data evaluation. Using AgenticMath, we construct a high-quality math dataset to fine-tune LLMs and enhance their mathematical reasoning ability. All agent prompts are provided in Appendix A.14.
@@ 3.2 Stage 1: Seed Problem Filtering
Problem Filtering Dimensions Complexity: Does it integrate multiple mathematical domains (e.g., algebra + geometry) or demand critical thinking? Information Value: Does it contain useful knowledge or reasoning opportunities? Can it help learners discover concepts, strategies, or patterns? Clarity & Precision: Is the question unambiguous, logically consistent, and free of errors? Poorly framed questions score lower.
@@ 3.3 Stage 2: Agentic Problem Generation
Problem Review Dimensions Clarity & Grammar: The question must be grammatically correct, precisely phrased, and easy to understand. It should avoid ambiguity in wording or phrasing. Logical Coherence & Completeness: All elements of the problem (e.g., given information, constraints, relationships, objectives) must be logically interconnected and sufficient. The problem should present a clear, sequential path for reasoning, without missing information required for the specified solution approach. Mathematical Validity & Solvability: The problem must be fundamentally a mathematics problem, with all its premises and conditions being mutually consistent and mathematically sound…
@@ 3.4 Stage 3: Solution Generation
$\mathcal{D}^{\text{sol}}_{\text{seed}}$ $=\{(q_{i}^{\text{seed}},\,a_{i}^{\text{sol}})\}_{i=1}^{N},$
$\mathcal{D}^{\text{sol}}_{\text{ref}}$ $=\{(q_{j}^{\text{ref}},\,a_{j}^{\text{sol}})\}_{j=1}^{K}.$
@@ 3.5 Stage 4: Synthetic Data Evaluation
In this stage, the scoring and curation framework from Stage 1 is extended to problem–solution pairs. The evaluation targets the synthetic set $\mathcal{D}^{\text{sol}}_{\text{ref}}=\{(q_{j}^{\text{ref}},a_{j}^{\text{sol}})\}_{j=1}^{K}$. Each pair is judged along three dimensions—clarity of the problem, correctness of the solution, and completeness of reasoning. Scores are further stabilized using the Score Transition Matrix and refined through $k$-NN consistency checks. To build a high-quality and diverse subset, a ranking-based selection is applied instead of a fixed threshold. Pairs are first sorted by quality and grouped into discrete score levels. We first group all samples by quality score (from 5 down to 0) and select groups in descending order. When the remaining quota falls within a group larger than needed, we rank samples inside that group using the long-tail diversity score. This strategy ensures that we always take the highest-quality data available while promoting diversity when selecting from oversized groups. This yields the curated dataset $\mathcal{D}_{\text{selected}}^{\text{sol}}=\{(q_{j}^{\text{ref}},a_{j}^{\text{sol}})\}_{j=1}^{L}$. The final training dataset combines this curated set with the seed-based solutions: $\mathcal{D}_{\text{final}}=\mathcal{D}_{\text{selected}}^{\text{sol}}\cup\mathcal{D}^{\text{sol}}_{\text{seed}}$, ensuring both rigor and diversity for downstream fine-tuning.
@@
DeepSeekMath-7B-MathFusion-7B 195K $58.2$ $79.5$ $40.3$ $69.1$ $25.5$ $27.0$ $49.9$
[…]
@@
[…]
@@ 4.1 Experimental Setup
We compare AgenticMath with state-of-the-art methods in two settings. For large-scale training, we include MetaMath (Yu et al., 2023), RFT (Yuan et al., 2023), DART-Math (Tong et al., 2024), MathScale (Tang et al., 2024), DeepSeekMath-7B-Instruct (sha, 2024), RefAug (Zhang et al., 2024c), MMIQC (Liu et al., 2025a), and WizardMath (Luo et al., 2023), all using 400K–2.3M samples. For small-scale, we evaluate 30K and 60K subsets: RefAug and MathFusion provide native 30K versions, while other baselines are randomly down-sampled to 60K from the large-scale dataset above.
@@ 4.3 Understanding AgenticMath: Ablations and Insights
Table 3: Ablation study on the contribution of different components of our method.
Method Samples AVG
Problem Rephrase 15K 31.4
   + Seed Filtering 15K 32.0 $\uparrow 0.6$
   + Problem Review–Revise 15K 33.0 $\uparrow 1.0$
   + Synthetic Data Evaluation 15K 33.2 $\uparrow 0.2$
We conduct ablation studies on Mistral-7B with a fixed 15K synthesized dataset to analyze the contribution of each AgenticMath component. As shown in Table 3, our method demonstrates consistent stage-wise improvements. Seed Filtering and the Review–Revise loop both yield clear performance gains, with iterative refinement playing a particularly important role. The final Synthetic Data Evaluation stage provides an additional boost, and the full method achieves the best overall performance. Additional ablation can be find in Appendix A.6.1A.6.2.
@@
Improved Performance under the Same Teacher Model.
To control for the influence of the teacher model, we evaluate AgenticMath against prior work using the same teacher model. As detailed in the appendix A.5, AgenticMath consistently outperforms previous methods under identical teacher supervision, indicating that the improvements arise from framework design rather than just stronger teachers.
5 CONCLUSION
Limitations
Despite the strong performance of AgenticMath on mathematical data synthesis, several limitations remain. First, our current study focuses primarily on text-based mathematical problems, and extending the proposed agentic method to multimodal settings (e.g., problems involving diagrams or visual reasoning) remains an important direction for future work. Second, due to computational and resource constraints, we do not conduct large-scale data synthesis beyond the explored sample sizes. Evaluating the scalability of AgenticMath under substantially larger data budgets is left for future investigation.
Appendix A Appendix
A.1 The Use of Large Language Models
A.2 Training and inference Details
[…]
During inference, we fix the sampling temperature to 0 to ensure deterministic outputs, and set the maximum generation length (max tokens) to 2048 for all models. We use a fixed random seed of 0 for reproducibility and set the number of inference trials to 1 for every evaluation. For our primary models, we adopt a standard Chain-of-Thought (CoT) prompting scheme. Specifically:
  • Training prompt: Question: {problem} Answer:
  • Evaluation prompt: Question: {problem} Answer: Let’s think step by step.
This prompt design follows common practice in mathematical reasoning tasks and encourages the model to generate explicit intermediate reasoning steps. For Mistral 7B and Llama 3 8B, we instead use the Alpaca instruction-following template during inference:
Below is an instruction that describes a task. Write a response that appropriately completes the request:
### Instruction:
{problem}
### Response:
We adopt this template because our preliminary experiments showed that Alpaca-style instructions consistently yield better reasoning quality on these two architectures compared with the CoT-style prompt. This observation is also aligned with the findings reported in MathFusion, where Alpaca-style prompting was similarly found to be more effective.
Table 5: 10K-sample comparison using the same teacher model.
Dataset # Samples MATH GSM8K College DM Olympiad Theorem AVG
MetaMath 10K $24.9$ $70.4$ $19.0$ $28.5$ $5.1$ $13.8$ $26.9$
DARTMath 10K $29.3$ $66.4$ $21.3$ $37.4$ $8.7$ $16.7$ $29.9$
ScaleQuest 10K $24.9$ $66.4$ $17.5$ $27.3$ $7.7$ $14.1$ $26.3$
[…]
Table 6: Sensitivity analysis of the revise threshold $\tau_{\text{rev}}$ using Llama3-8B.
$\tau_{\text{rev}}$ Samples MATH GSM8K College DM Olympiad Theorem AVG
4.5 30K $36.8$ $78.4$ $29.6$ $40.3$ $11.4$ $20.4$ $36.2$
4.0 30K $36.6$ $77.5$ $28.2$ $43.1$ $11.5$ $20.0$ $36.2$
3.5 30K $37.8$ $77.4$ $27.4$ $41.0$ $10.3$ $20.0$ $35.7$
A.3 Evaluation Benchmarks
In-Domain: (i) GSM8K (Cobbe et al., 2021), consisting of grade-school arithmetic word problems that are relatively simple. (ii) MATH (Hendrycks et al., 2021), a large-scale dataset of competition-level problems that are significantly more challenging.
Out-of-Domain (OOD): (i) CollegeMath (Tang et al., 2024), with 2,818 college-level problems drawn from nine textbooks across seven domains (e.g., linear algebra, differential equations), designed to test generalization to complex mathematics. (ii) DeepMind-Mathematics (Saxton et al., 2019), a collection of 1,000 problems covering a national school curriculum (up to age 16), assessing basic reasoning across varied types. (iii) OlympiadBench-Math (He et al., 2024), providing 675 Olympiad-level problems (English text-only subset) targeting the most challenging reasoning tasks. (iv) TheoremQA (Chen et al., 2023), consisting of 800 problems that require applying mathematical theorems across mathematics, physics, and engineering, testing theoretical reasoning in STEM.
A.4 Larger Training Samples Yield Stronger Reasoning Performance.
@@ A.3 Evaluation Benchmarks
A.5 Fairness Concerns Regarding Teacher Models
For further strengthen the fairness discussion, we conduct an additional 10K-sample controlled comparison across MetaMath, DARTMath, ScaleQuest, and our AgenticMath, all trained under the same SFT configuration (Mistral-7B) and, where applicable, using the same teacher model (GPT-4o-mini, 2024-07-18) for solution generation. As shown in Table 5, AgenticMath achieves the highest average performance among all methods. This controlled experiment confirms that the observed improvements do not arise from using a stronger teacher model—since all datasets share the same teacher—but rather from the design of our synthesis method itself.
A.6 Additional Ablations
A.6.1 Sensitivity Analysis on the Revise Threshold
To further examine the impact of the revise threshold, we conduct a sensitivity study with $\tau_{\text{rev}}\in\{3.5,\,4.0,\,4.5\}$ using Llama3-8B under a fixed 30K SFT setting. As shown in Table 6, the settings $\tau_{\text{rev}}=4.0$ and $4.5$ produce highly consistent results, demonstrating that the review–revise mechanism remains stable across reasonable threshold choices. In contrast, $\tau_{\text{rev}}=3.5$ yields lower overall performance, which is expected since a looser threshold admits more low-quality candidates into subsequent stages, ultimately weakening the final dataset quality.
A.6.2 Analysis of the Three Review–Revise Iterations.
To further address the reviewer’s question regarding the choice of three review–revise iterations, we evaluate the quality of the problems that pass each round using GPT-4o-mini. Specifically, we compute the average complexity, information value, and clarity scores for all accepted problems after each iteration.
Table 7: Quality metrics across three review–revise iterations.
Metric Round 1 Round 2 Round 3
Complexity $3.86$ $3.93$ $3.92$
Information Value $3.96$ $4.03$ $4.02$
Clarity $4.35$ $4.44$ $4.45$
Avg 4.06 4.13 4.13
As shown in Table 7, the first two review–revise iterations produce consistent improvements across all metrics. By the third iteration, however, the gains largely stabilize, indicating diminishing returns. This analysis supports our design choice of using three iterations: it captures most of the quality improvements while avoiding unnecessary computation beyond the point of saturation.
Table 8: Score distribution for GSM8K and MATH seed datasets.
Dataset Score=0 Score=1 Score=2 Score=3 Score=4 Score=5 Score $\geq 3$
GSM8K 823 3522 1849 1188 91 0 1279
MATH 69 825 884 2053 3415 254 5722
A.7 Additional Analysis of Synthetic Data Quality and Characteristics
To provide further quantitative evidence of the quality and semantic composition of the synthesized data, we conduct a post-hoc analysis using GPT-4o-mini as an external evaluator. Specifically, we (i) assign quality scores to the refined problems and (ii) classify each problem into standard mathematical topics.
A.7.1 Quality Score Distribution
Table 9 reports the quality distribution assigned by GPT-4o-mini over the 18,679 refined problems. The majority of synthesized questions receive a score of $\geq 4$, indicating strong clarity, coherent reasoning, and non-trivial complexity across the dataset.
Table 9: Quality score distribution of 18,679 refined problems, evaluated by GPT-4o-mini.
Quality Score Count Percentage
1 307 1.64%
2 1175 6.29%
3 5053 27.05%
4 12142 65.00%
5 2 0.01%
A.7.2 Topic Distribution of the Final 15K Dataset
GPT-4o-mini is further used to classify each problem in the final 15,000-sample dataset into standard mathematical domains. The resulting topic distribution is shown in Table 10. The distribution demonstrates broad semantic coverage across major mathematical disciplines, with strong representation in combinatorics, geometry, algebra, and number theory.
Table 10: Topic distribution of the final 15K synthetic dataset (classified by GPT-4o-mini).
Topic Count Percentage
Counting & Probability 3705 24.70%
Geometry 3326 22.17%
Algebra 2446 16.31%
Number Theory 1475 9.83%
Calculus 1470 9.80%
Precalculus 1456 9.71%
Intermediate Algebra 907 6.05%
Prealgebra 123 0.82%
Linear Algebra 71 0.47%
Others 21 0.14%
A.8 Detailed Statistics of the Data Generation Method
To provide a clearer understanding of the robustness and transparency of our data generation method, we report detailed statistics for all major stages, including seed scoring, rephrase expansion, the multi-round review–revise refinement process, and the final data evaluation. These analyses illustrate how the method progressively improves problem diversity, clarity, and complexity, which are key for enhancing downstream mathematical reasoning ability.
A.8.1 Seed Data Scoring
We first evaluate all raw seed questions using our label-free scoring mechanism. A total of 7,001 questions satisfy the filtering threshold (score $\geq 3$). The full distribution is shown in Table 8. This stage ensures that only seed questions with sufficient structural soundness and baseline complexity are used for further synthesis.
A.8.2 Rephrase Expansion
To enhance problem complexity and diversity while preserving core semantics, each filtered seed question is rephrased six times. All 42,006 synthesized candidates proceed to the review–revise process. This expands the problem pool as follows:
Table 11: Rephrase expansion of filtered seed questions.
Stage Count Calculation Total
Rephrase Expansion $(1279+5722)\times 6$ 42,006
A.8.3 Review–Revise Loop
Our three-round review–revise process progressively improves clarity, logical correctness, and mathematical validity. Across rounds, vague or low-quality questions are removed, while clearer and more coherent problems are retained. Table 12 summarizes the filtering behavior across rounds.
Table 12: Statistics of the three-round review–revise refinement.
Round Total Inputs Passed Pass Rate
1 42,006 7,438 17.71%
2 34,568 6,526 18.88%
3 28,042 4,718 16.83%
All Rounds 18,682
A.8.4 Synthetic Data Quality Distribution
After refinement, 18,679 high-quality synthetic problems remain. The fact that 65% of the questions are assigned a score of 4, with another 27% scoring 3, demonstrates that the majority of synthesized problems exhibit strong clarity, coherent reasoning, and meaningful complexity. Their quality distribution (evaluated by GPT-4o-mini) is shown in Table 13.
Table 13: Quality score distribution of 18,679 refined synthetic problems.
Score Count Percentage
1 307 1.64%
2 1175 6.29%
3 5053 27.05%
4 12,142 65.00%
5 2 0.01%
A.8.5 Final Dataset Construction
To construct the final 15K dataset used in our experiments, we jointly rank all refined samples using a combination of the quality score and the long-tail diversity score. This ranking procedure prioritizes both overall quality and distributional diversity. The top 15,000 problems from this ranked list form the final synthetic dataset.
A.9 Clarifying the Novelty of AgenticMath
Our approach introduces several key components that are under-explored or not systematically addressed in existing math data-generation methods. The main distinctions are summarized as follows:
  • Seed Question Filtering. Most prior approaches synthesize new problems from the entire seed set without assessing seed quality. In contrast, we explicitly score and filter seed questions before synthesis, which improves the quality of generated problems, avoids unnecessary synthesis from weak or ambiguous seeds, and leads to stronger downstream performance under the same data budget.
  • Iterative Review–Revise Loop for Problem Quality. Existing methods often emphasize increasing the volume or diversity of generated problems while overlooking intrinsic quality. As a result, some synthesized problems may contain unclear statements, logical inconsistencies, or may not admit valid solutions. Our dedicated review–revise loop directly addresses these issues by iteratively refining clarity, structure, and mathematical solvability before problems proceed to later stages.
  • Joint Evaluation of Problems and Solutions. Because synthetic problems lack ground-truth labels, many prior methods do not systematically verify the correctness of generated solutions. Our final evaluation stage explicitly scores both the generated problem and its corresponding solution, filtering out samples with incorrect reasoning or inconsistent final answers. This joint evaluation plays a critical role in ensuring overall data reliability.
  • Independence from Seed Solutions. Unlike methods that rely on ground-truth seed solutions to guide synthesis or estimate problem difficulty, our method operates solely on the seed question text. This removes the dependency on labeled solutions, reduces annotation requirements, and enables broader applicability in scenarios where ground-truth solutions are unavailable.
Overall, these components make our method substantially more robust and systematic compared to existing approaches. They collectively contribute to the strong empirical gains observed in our experiments and highlight the benefits of explicitly modeling data quality throughout the synthesis process.
A.10 Comparison Between Synthetic Questions and Seed Questions
To examine the quality of purely synthetic questions relative to the original seed questions, we conduct a controlled comparison between the original 15K seed dataset and a 15K dataset generated purely by AgenticMath.
To ensure fairness, solutions for both datasets are rewritten using the same teacher model (GPT-4o-mini), and all models are trained under identical SFT settings.
Table 14 shows that the dataset generated by AgenticMath consistently outperforms the original seed dataset across all benchmarks. In particular, AgenticMath achieves a substantially higher overall average score (33.2 vs. 29.3). More importantly, the gains are especially pronounced on out-of-domain benchmarks, including CollegeMath, DeepMind Mathematics (DM), OlympiadBench, and TheoremQA.
These results demonstrate that the synthetic questions produced by AgenticMath not only match but frequently exceed the quality of the original seed questions. The stronger out-of-domain performance indicates that AgenticMath generates problems with richer structures, more diverse reasoning patterns, and improved transferability. Overall, this comparison confirms that AgenticMath does not merely replicate the seed distribution, but instead synthesizes novel mathematical problems that generalize more effectively to unseen domains.
A.11 Related work
Data Selection.
Data selection has been shown to improve instruction tuning, achieving performance comparable to full training datasets (Xia et al., 2022; Lu et al., 2023; Xia et al., 2023; Li et al., 2024c; Xia et al., 2024; Pang et al., 2024; Zhang et al., 2024b; Long et al., 2024; Li et al., 2025a; Wang et al., 2025; Xu et al., 2025; Pang et al., 2025; Li et al., 2025b). Specifically, LESS (Pang et al., 2024) proposes introduces a low-rank gradient similarity–based selection method. DS2 (Pang et al., 2024) further refines data quality by rating and curating samples with LLM. In this work, we integrate LLM-based selection into data generation, allowing synthesis conditioned solely on high-quality questions.
Table 14: Comparison between original seed questions and purely synthetic questions generated by AgenticMath.
Method # Samples MATH GSM8K College DM Olympiad Theorem AVG
Mistral-7B: GSM8K + MATH 15K $28.6$ $71.1$ $20.3$ $33.4$ $6.8$ $15.8$ $29.3$